OpenStudy (anonymous):

Which of the following is greatest? seven twenty-fifths zero point three three nineteen one hundredths eleven-fiftieths

OpenStudy (anonymous):

0.33

OpenStudy (anonymous):

So here's what you have: 7/25 0.33 0.19 11/50 The easiest way is to convert them all to decimals. 7/25 = 0.28 11/50 = 0.22 (you can use a calculator or google for this) Then you can see which is the greatest.

OpenStudy (anonymous):

0.28 0.33 0.19 0.22 0.33 is the greatest. Do you get it?